You want to make a 5-year investment on bonds, and if you buy bonds with longer period of maturity, they will be sold at the prevailing market price at the end of the fifth year. All coupons will be reinvested. You forecast that the rates are going to change after you purchase the bonds. Now you have three bonds with same initial YTM to consider: A. 5-year maturity, 6% coupon paid annually, YTM=9% B. 10-year maturity, 8% coupon paid annually, YTM=9% C. 15-year maturity, 4% coupon paid annually, YTM=9% To pick the best one, you want to calculate the following:

1) What are the current prices of three bonds based on the initial YTM given? [2 points]

2) If the coupon payments are reinvested at the prevailing market rate (7% for the coming 5 years), what are the future values (at the end of the fifth year) of all five coupon payments for three bonds respectively? [3 points]

3) What are the selling prices of three bonds at the end of the fifth year, given that the prevailing market rate is 8% at that time? (For bond A, use its par value as the final selling price, as it will mature in 5 years). [2 points]

4) What are the HPRs for the three bonds? [2 points]

5) What’s your recommendation? [1 point]
Answer #1

ANSWER

1)

1,000.00 6.00% 60.00 1 Bond par value 2 Coupon rate 3 Number of compounding periods per year 4 = 1*2/3 Interest per period (P

1,000.00 8.00% 1 Bond par value 2 Coupon rate 3 Number of compounding periods per year 4 = 1*2/3 Interest per period (PMT) 5

1,000.00 4.00% 1 Bond par value 2 Coupon rate 3 Number of compounding periods per year 4 = 1*2/3 Interest per period (PMT) 5

2)

Rate 7% = 0.07 Nper 5 = 5 Pmt -1000*6% = -60 Pv| = number Type = number = 345.0443406 Returns the future value of an investme

Future value of Bond A=$345.04

Rate 7% Nper Pmt -1000*8% Pv Type = 0.07 = 5 = -80 = number = number * = 460.0591208 Returns the future value of an investmen

Future value of Bond B=$460.06

* = 0.07 Rate 7% Nper 5 Pmt -1000*4% PV * = -40 = number = number Type = 230.0295604 Returns the future value of an investmen

Future value of the bond C =$230.03

3)
Selling price of bond A=$1000

Bond B coupon rate and market rate are same therefore selling price of bond B is $1000

1,000.00 4.00% 40.00 1 Bond par value 2 Coupon rate 3 Number of compounding periods per year 4 = 1*2/3 Interest per period (P

Selling price of the bond C is $731.60

4)

1,000.00 6.00% 1 Bond sale price after 5 year 2 Coupon rate 3 Number of compounding periods per year 1*2/3 Interest per perio

HPR of bond A is 9%

1,000.00 8.00% 1 Bond sale price after 5 year 2 Coupon rate 3 Number of compounding periods per year 1*2/3 Interest per perio

HPR of bond B is 9.68%

731.60 4.00% 40.00 1 Bond sale price after 5 year 2 Coupon rate 3 Number of compounding periods per year 1*2/3 Interest per p

HPR of Bond C is 10.37%

5)

The Purchase price of the bond C is less and the holding period return is high for bond C. Hence it is recommended to purchase Bond C.

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************